gelphielvr:

(algebra 1) linear relationships, slope, table hii I need help getting started & possibly solving this question / posted in the replies

gelphielvr:

1 attachment
ThanosQC:

To find the slope of this line, use slope formula: (\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1}}\]). \[\frac{ 18-28 }{ -5+9 }=\frac{ -10 }{ 4 }=-\frac{ 5 }{ 2 }\]

gelphielvr:

@thanosqc wrote:
To find the slope of this line, use slope formula: (\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1}}\]). \[\frac{ 18-28 }{ -5+9 }=\frac{ -10 }{ 4 }=-\frac{ 5 }{ 2 }\]
thank you Thanos - I always get confused with the y over x